# Attention Pruning: Automated Fairness Repair of Language Models via Surrogate Simulated Annealing

### v2

@@ Abstract.
This paper explores pruning attention heads as a post-processing bias mitigation method for large language models (LLMs). LLMs have been expanding into sensitive social contexts and socio-economic decision-making where fairness concerns become especially crucial. Since LLMs develop their decision-making patterns by training on massive datasets of human-generated content, they naturally encode and perpetuate societal biases. While modifying training datasets and algorithms is prohibitive, post-processing techniques—such as pruning attention heads in pre-trained LLMs—can provide feasible and effective approaches to improve fairness. However, identifying the optimal subset of parameters to prune presents a combinatorial challenge within the immense parameter space of LLMs, requiring efficient solutions that efficiently balance competing objectives across the frontiers of model fairness and utility.
@@
conference: 48th International Conference on Software Engineering (ICSE); 12-18 April, 2026; Rio de Janeiro, Brazilcopyright: none
@@ 1. Introduction
Recent advances in large language models (LLMs) have enabled near-human performance on various tasks (OpenAI, 2024) and have catalyzed their widespread adoption in real-world use cases. However, despite their impressive performance, prior work has shown that these models often inherit biases from training data (Hada et al., 2023). To make matters worse, such biases can produce uneven performance across different social groups in society (Gupta et al., 2023; Li et al., 2020; Nagireddy et al., 2024; Guo et al., 2024b). Bias in language models is particularly troubling given its potential for significant negative societal impacts (United Nations Educational, Scientific and Cultural Organization, 2024; Milne, 2024; Gallegos et al., 2024). Consequently, equipping LLM-based software developers to mitigate bias in socially critical AI systems is essential to promote inclusivity and equality.
@@
Significant research within the software engineering (SE) community has focused on addressing discrimination in data-driven software systems (Galhotra et al., 2017; Aggarwal et al., 2019; Udeshi et al., 2018; Zhang et al., 2020; Chakraborty et al., 2020). Fairness has emerged as a crucial meta-functional property, requiring analysis beyond functional correctness and evaluation metrics that go beyond prediction accuracy (Brun and Meliou, 2018). Consequently, researchers have developed various testing (Galhotra et al., 2017; Aggarwal et al., 2019; Zhang et al., 2020), debugging (Biswas and Rajan, 2021; Monjezi et al., 2023; Yu et al., 2025), and repairing (Dasu et al., 2024; Peng et al., 2023; Chen et al., 2022; Li et al., 2023; Gao et al., 2022) approaches to address fairness issues in data-driven software systems. However, repairing unfairness in the frontier AI models has been understudied within the SE community.
@@
Various techniques have been proposed to mitigate social biases in language models within the AI community. Some past works aim to find a subset of weights containing bias and then modify only these weights to improve fairness (Meissner et al., 2022; Le Bras et al., 2020). Others tried to unlearn the undesired properties by modifying the representations (Li et al., 2024). Recently, Zayed et al. (Zayed et al., 2024) presented a pruning strategy to mitigate unfairness in LLMs that outperformed prior works. Despite all these efforts, improving fairness in LLMs while balancing the inevitable trade-off in model utility has remained a challenge.
@@
Inspired by search-based program repair techniques  (Qi et al., 2014; Forrest et al., 2009; Ruan et al., 2024), we propose a randomized search algorithm to semantically repair unfairness in LLMs without degrading their functional utilities. Our key observation is that attention heads in LLMs may disparately contribute to unfairness while minimally affecting utility. Hence, pruning an ideal set of attention heads as a post-processing mitigation strategy can improve fairness with a negligible impact on utility. However, pruning is a combinatorial search problem that requires an exhaustive search over all possible subsets of attention heads and poses a significant computation challenge in LLMs with over a thousand attention heads. Prior work on fairness-aware attention head pruning (Zayed et al., 2024) avoids the combinatorial problem by only analyzing the effect each attention head has on fairness. Such an approach fails to account for different combinations of attention heads and their complex nonlinear relationships.
@@
We propose a randomized heuristic search in the class of simulated annealing (SA) algorithms to explore the search space efficiently. However, due to the high inference time of billion-parameter LLMs, a direct application of SA is infeasible, as exploring even a single state takes several minutes. We make a key observation that the effect of pruning a subset of attention heads on fairness/utility follows a pattern and can be learned using surrogate functions via deep neural networks (DNNs). This novel use of surrogate functions scales the SA search to find an ideal subset of attention heads in the surrogate models rather than the LLMs.
@@
We implement and evaluate Attention Pruning (AP), an automated fairness repair strategy for LLMs using SA. AP identifies an ideal subset of attention heads to prune such that the fairness of LLMs improves while ensuring minimal loss in model utility. It uses surrogate DNNs during the SA search to explore the combinatorial search space of billion-parameter LLMs efficiently. We investigate the following research questions with 6 LLMs, focusing on reducing gender bias from the HolisticBias (Smith et al., 2022) dataset:
@@
RQ3) Design considerations of Attention Pruning. We examine how different hyperparameters can be tuned to have fine-grained control over the fairness-utility trade-off. We identify one hyperparameter that can effectively explore the Pareto frontier of fairnesss-utility in LLMs and balance their trade-off.
@@
    Approximating the Fairness and Utility of LLMs We show that the fairness and utility of pruned LLMs can be accurately approximated using small surrogate deep neural networks,
@@
    Surrogate Simulated Annealing We propose an efficient discrete optimization algorithm using a surrogate simulated annealing that handles billion-parameter LLMs and outperforms the state-of-the-art bias mitigators
@@
    Experimental Evaluation We present AP and evaluate its effectiveness and efficiencies in reducing social biases in various LLMs.
@@ 2.2. Bias in Causal Language Modeling
Bias in LLMs for causal language modeling can be broadly categorized into two categories (Guo et al., 2024a): intrinsic bias and extrinsic bias. Intrinsic biases arise from the training dataset, learning algorithms, or the model architecture itself (Sun et al., 2019). LLMs trained on vast text sequences from the Internet tend to inherit the biases present in them. For example, LLMs reinforce gender stereotypes by associating the word “doctor” with men and “nurse” with women (Bolukbasi et al., 2016). The side effects of such biases are amplified when they are employed in sensitive socio-economic applications.
Extrinsic biases in LLMs manifest in specific real-world applications. As a result, the metrics and benchmarks that have been proposed to evaluate extrinsic bias are task-specific. For example, Sap et al. (2019) shows that tweets written in African American English (AAE) by self-identified African-Americans are two times more likely to be labeled as offensive than other tweets. Other benchmarks (Rudinger et al., 2018) contains a set of Winograd-schema style sentences for coreference resolution focused on gender bias. For example, given the sentence “the doctor asked the nurse to help her because she was busy”, models often erroneously correlate “her” with “nurse” rather than with “doctor”.
@@ 2.3. Notions of Model Utility via Perplexity
A lower perplexity score implies that model $\Theta$ has been trained well to estimate the real-world probability distribution. Informally, a sequence with a low perplexity score implies that the model is less “surprised” by a sequence of tokens. In our work, we measure the perplexity on the WikiText-2 (Merity et al., 2016) dataset.
@@ 2.4. Notions of Model Fairness
In our work, we use the HolisticBias (Smith et al., 2022) metric that measures a form of extrinsic bias. The HolisticBias metric builds on BOLD (Dhamala et al., 2021) and contains a dataset of 566,000 prompts that are categorized into 13 social biases (e.g., Gender, Sexual Orientation, Race and Ethnicity, Nationality, etc.). Bias is quantified by prompting the LLMs with sentences from the different groups and examining the toxicity of the generated text. The toxicity of a text determines how disrespectful, unpleasant, or harmful the language is. The toxicity is computed using a BERT (Devlin et al., 2019) model that is pre-trained on a toxic comment classification dataset [^1] https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge. The BERT model outputs a score in $[0,1]$ where a higher score is assigned to more toxic text.
To measure the bias of the LLM, we adopt the procedure outlined in prior works (Zayed et al., 2024; Dhamala et al., 2021). Each social bias, or group, in the HolisticBias dataset consists of a set of subgroups $G$. For example, the gender group contains subgroups such as binary, queer, transgender, etc. The bias of the LLM $\Theta$ on a set of subgroups $G$ is
$\text{bias}_{\Theta}(G)=\sum_{g\in G}\left|T_{G}-T_{g}\right|$
$\text{where}\quad T_{g}=\frac{1}{|D_{g}|}\sum_{x\in D_{g}}\text{tox}_{\Theta}(x)$
$\text{and}\quad T_{G}=\frac{1}{|G|}\sum_{g\in G}T_{g}$
@@ 3.1. Multi-Head Attention in Transformers
Most modern LLMs for causal language modeling are based on the Transformer architecture proposed by Vaswani et al. (2017). On a high level, the Transformer is a neural network that comprises $n$ blocks stacked on top of each other, where each block has a Multi-Head Attention (MHA) layer followed by a fully connected layer. The MHA layer utilizes scaled-dot product attention, which operates on a set of queries $Q$, keys $K$, and values $V$, which come from the previous layer’s output. The query, key, and values are computed from the input text embeddings for the first block. Scaled-dot product attention is computed as
@@ 3.2. Attention Head Pruning for Fairness
Pruning in LLMs refers to the practice of “pruning” or “dropping” parts of the LLM. Pruning strategies can be broadly categorized into two categories: unstructured and structured pruning. Unstructured pruning refers to the practice of pruning at the neuron level, where a subset of neurons are pruned. On the other hand, structured pruning strategies remove entire structural components, such as attention heads or layers, from the LLM. Attention head pruning is a structured pruning strategy that prunes one or more attention heads from multi-head attention in a Transformer block. For example, Figure 1(b) shows multi-head attention with the second and fourth heads pruned. These heads are not concatenated before multiplying with the output project matrix.
@@
Multiple works have been proposed (Michel et al., 2019; Voita et al., 2019; Fan et al., 2021; Han et al., 2016; Behnke and Heafield, 2020) to prune redundant attention heads in LLMs that have a negligible impact on utility. However, none of these methods account for the model’s fairness and only focus on pruning heads while minimizing loss to model utility. We define the problem of pruning attention heads to improve fairness while minimizing loss to model utility as follows
@@
Recently, Zayed et al. (2024) proposed the fairness-aware structured pruning (FASP) algorithm that prunes attention heads while accounting for both fairness and model utility. FASP quantifies the contribution of an attention head by measuring the change in bias and perplexity of LLM before and after pruning it. Formally, for a model with $N$ attention heads, the impact of each head $h$ on a social group represented by set $G$ is estimated as
@@
As the authors concede, a major shortcoming of FASP is that it fails to account for the complex non-linear relationship of attention heads across multiple Transformer blocks in an LLM. Specifically, the effect of dropping heads one at a time is not the same as the effect of dropping multiple heads simultaneously. For example, if dropping the first and second heads decreases bias while not reducing perplexity, dropping both of them together may not decrease bias, and it might worsen perplexity. Nevertheless, they argue that FASP is a first step toward fairness-aware attention head pruning.
@@ 4.1. Discrete Optimization with Simulated Annealing
where the set of vertices $V$ consists of all bitstrings of length $n$
@@
The cooling schedule determines how the temperature changes over time. In SA, we start with a high temperature and anneal it over time. This encourages exploration in the early stages of the SA search and exploitation towards the end. We adopt the logarithmic cooling schedule (Hajek, 1988). The temperature $T_{i}$ at the $i^{th}$ iteration is defined as
@@
where $T_{0}$ is the initial temperature. We utilize the temperature initialization algorithm proposed by Ben-Ameur (2004) to set $T_{0}$.
@@ 5. Experiments
Datasets and Models. We primarily focus on reducing gender bias from the HolisticBias metric in LLMs. We consider 6 LLMs for our analysis: Distilgpt-2 (Sanh et al., 2019), GPT-2 (Radford et al., 2019), GPT-Neo-125M (Black et al., 2021), GPT-Neo-1.3B (Black et al., 2021), GPT-J-6B (Wang and Komatsuzaki, 2021), Llama-2-7B (Touvron et al., 2023). Our choice of social bias and LLMs mimics the state-of-the-art FASP pruning strategy setup for a fair comparison. The LLMs chosen cover a diverse set of parameter sizes, from 82M (Distilgpt-2) all the way to 7B (Llama-2-7B).
The architecture for the surrogate DNNs is shown in Table 2. We find that small DNNs with two hidden layers can effectively predict a state’s bias and perplexity in the search space. The size of the input layer is the number of attention heads in the LLM. The output layer has one neuron. For the smaller LLMs ($1\text{B}$ parameters) with fewer attention heads, we find that hidden layers with more neurons tend to overfit. Thus, we opt to use small hidden layers with 64 and 32 neurons. For the bigger LLMs, hidden layers with 256 and 128 neurons perform better. The size of the datasets used to train the surrogate DNNs is highlighted in Table 2. We use a train-validation split of 95%-5%. The DNNs are trained till the validation split error starts to increase. During the dataset creation in Algorithm 1, we set the upper bound on the number of attention heads dropped $n_{u}$ to up to $10\%-20\%$ of the total attention heads in the LLM. We set $\eta_{bias}\in[0.1,0.2]$ and $\eta_{ppl}=1$. Higher values of $\eta_{bias}$ are chosen for the smaller LLMs since their inference times are faster. We use the entire WikiText-2 validation split as it only contains $2,461$ text sequences. Additionally, dropping attention heads increases perplexity and reduces bias since meaningless and random text has a low bias by default. Therefore, we need the perplexity DNN $\theta_{ppl}$ to be as accurate as possible to ensure that the SA search returns states with low bias and perplexity. To ensure high data quality for the perplexity samples, we used the entire WikiText-2 validation split. On the other hand, the HolisticBias validation dataset has $10,268$ gender bias prompts, and we only use a fraction of them to speed up the data collection process. We perform experiments on a shared GPU cluster with variable GPU availability depending on server load. We note that the dataset collection process is offline. Instead of setting a time limit for Algorithm 1, we collect at least 25,000 samples to train the DNNs. Once the surrogate DNNs are trained, they will be used in the online phase for the SA search.
@@
Baseline and Considerations. We compare our approach with four pruning strategies. The first is FASP (Zayed et al., 2024), which is the state-of-the-art fairness-aware pruning algorithm. The other two strategies prune heads based on weight magnitude (Han et al., 2016) and gradient magnitude (Michel et al., 2019). On a high level, these methods rank the heads based on how large the weights of the heads or gradients are with respect to a loss function after masking the heads. Heads with low weight and gradient magnitude are deemed unimportant. These methods are state-of-the-art general-purpose pruning algorithms. The fourth approach is random pruning, which serves as an ablation study.
@@
where $\mathcal{P}$ is the bias dataset, $p_{max}$ is the threshold for maximum perplexity, and $\mathcal{P^{\prime}}$ is the updated dataset after clamping $\mathcal{P}$, and $\sigma$ is the standard deviation threshold. We set $\sigma\leq 10$. After clamping, we scale down the dataset by $p_{max}$ to ensure that all values lie in $[0,1]$. We observe that this approach improves the performance of the DNNs during the search as they assign the value $p_{max}$ to states that degrade the model’s utility beyond acceptable levels.

### v3

@@ Abstract.
This paper explores pruning attention heads as a post-processing bias mitigation method for large language models (LLMs). LLMs have been expanding into sensitive social contexts and socio-economic decision-making where fairness concerns become especially crucial. Since LLMs develop their decision-making patterns by training on massive datasets of human-generated content, they naturally encode and perpetuate societal biases. While modifying training datasets and algorithms is prohibitive, post-processing techniques—such as pruning attention heads in pre-trained LLMs—can provide feasible and effective approaches to improve fairness. However, identifying the optimal subset of parameters to prune presents a combinatorial challenge within the immense parameter space of LLMs, requiring efficient solutions that balance competing objectives across the frontiers of model fairness and utility.
@@
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773115isbn: 979-8-4007-2025-3/26/04copyright: none
@@ 1. Introduction
Recent advances in large language models (LLMs) have enabled near-human performance on various tasks (openai2024gpt4technicalreport) and have catalyzed their widespread adoption in real-world use cases. However, despite their impressive performance, prior work has shown that these models often inherit biases from training data (hada-etal-2023-fifty). To make matters worse, such biases can produce uneven performance across different social groups in society (gupta2023calm; li-etal-2020-unqovering; nagireddy2024socialstigmaqa; guo2024hey). Bias in language models is particularly troubling given its potential for significant negative societal impacts (UNESCO2024; Milne2024; gallegos2024bias). Consequently, equipping LLM-based software developers to mitigate bias in socially critical AI systems is essential to promote inclusivity and equality.
@@
Significant research within the software engineering (SE) community has focused on addressing discrimination in data-driven software systems (FairnessTesting; 10.1145/3338906.3338937; udeshi2018automated; ADF; chakraborty2020fairway). Fairness has emerged as a crucial meta-functional property, requiring analysis beyond functional correctness and evaluation metrics that go beyond prediction accuracy (brun2018software). Consequently, researchers have developed various testing (FairnessTesting; 10.1145/3338906.3338937; ADF), debugging (10.1145/3468264.3468536; 10.1109/ICSE48619.2023.00136; yu2024fairlay), and repairing (10.1145/3650212.3680380; 10.1109/TSE.2022.3220713; chen2022maat; 10.1145/3617168; 10.1145/3510003.3510087) approaches to address fairness issues in data-driven software systems. However, repairing unfairness in the frontier AI models has been understudied within the SE community.
@@
Various techniques have been proposed to mitigate social biases in language models within the AI community. Some past works aim to find a subset of weights containing bias and then modify only these weights to improve fairness (meissner-etal-2022-debiasing; le2020adversarial). Others tried to unlearn the undesired properties by modifying the representations (li2024wmdp). Recently, Zayed et al. (zayed2024fairness) presented a pruning strategy to mitigate unfairness in LLMs that outperformed prior works. Despite all these efforts, improving fairness in LLMs while balancing the inevitable trade-off in model utility has remained a challenge.
@@
Inspired by search-based program repair techniques  (10.1145/2568225.2568254; 10.1145/1569901.1570031; ruan2024evolutionary), we propose a randomized search algorithm to semantically repair unfairness in LLMs without degrading their functional utilities. Our key observation is that attention heads in LLMs may disparately contribute to unfairness while minimally affecting utility. Hence, pruning an ideal set of attention heads as a post-processing mitigation strategy can improve fairness with a negligible impact on utility. However, pruning is a combinatorial search problem that requires an exhaustive search over all possible subsets of attention heads and poses a significant computation challenge in LLMs with over a thousand attention heads. Prior work on fairness-aware attention head pruning (zayed2024fairness) avoids the combinatorial problem by only analyzing the effect each attention head has on fairness. Such an approach fails to account for different combinations of attention heads and their complex nonlinear relationships.
@@
We propose a randomized heuristic search in the class of simulated annealing (SA) algorithms to explore the search space efficiently. However, due to the high inference time of billion-parameter LLMs, a direct application of SA is infeasible, as exploring even a single state takes several minutes. We make a key observation that the effect of pruning a subset of attention heads on fairness/utility follows a pattern that can be learned using surrogate functions via deep neural networks (DNNs). This novel use of surrogate functions scales the SA search to find an ideal subset of attention heads using the surrogate models rather than the LLMs.
@@
We implement and evaluate Attention Pruning (AP), an automated fairness repair strategy for LLMs using SA. AP identifies an ideal subset of attention heads to prune such that the fairness of LLMs improves while ensuring minimal loss in model utility. It uses surrogate DNNs during the SA search to explore the combinatorial search space of billion-parameter LLMs efficiently. We investigate the following research questions with 6 LLMs, focusing on reducing gender bias from the HolisticBias (smith2022imsorryhearthat) dataset:
@@
RQ3) Design considerations of Attention Pruning. We examine how different hyperparameters can be tuned to have fine-grained control over the fairness-utility trade-off. We identify one hyperparameter that can effectively explore the Pareto frontier of fairness-utility in LLMs and balance their trade-off.
@@
    Approximating the Fairness and Utility of LLMs. We show that the fairness and utility of pruned LLMs can be accurately approximated using small surrogate deep neural networks.
@@
    Surrogate Simulated Annealing. We propose an efficient discrete optimization algorithm using surrogate simulated annealing that handles billion-parameter LLMs and outperforms the state-of-the-art bias mitigators.
@@
    Experimental Evaluation. We present AP and evaluate its effectiveness in reducing social biases in various LLMs.
@@ 2.2. Bias in Causal Language Modeling
Bias in LLMs for causal language modeling can be broadly categorized into two categories (guo2024biaslargelanguagemodels): intrinsic bias and extrinsic bias. Intrinsic biases arise from the training dataset, learning algorithms, or the model architecture itself (sun-etal-2019-mitigating). LLMs trained on vast text sequences from the Internet tend to inherit the biases present in them. For example, LLMs reinforce gender stereotypes by associating the word “doctor” with men and “nurse” with women (bolukbasi2016mancomputerprogrammerwoman). The side effects of such biases are amplified when they are employed in sensitive socio-economic applications.
Extrinsic biases in LLMs manifest in specific real-world applications. As a result, the metrics and benchmarks that have been proposed to evaluate extrinsic bias are task-specific. For example, sap-etal-2019-risk shows that tweets written in African American English (AAE) by self-identified African-Americans are two times more likely to be labeled as offensive than other tweets. Other benchmarks (rudinger2018genderbiascoreferenceresolution) contains a set of Winograd-schema style sentences for coreference resolution focused on gender bias. For example, given the sentence “the doctor asked the nurse to help her because she was busy”, models often erroneously correlate “her” with “nurse” rather than with “doctor”.
@@ 2.3. Notions of Model Utility via Perplexity
A lower perplexity score implies that model $\Theta$ has been trained well to estimate the real-world probability distribution. Informally, a sequence with a low perplexity score implies that the model is less “surprised” by a sequence of tokens. In our work, we measure the perplexity on the WikiText-2 (merity2016pointersentinelmixturemodels) dataset.
@@ 2.4. Notions of Model Fairness
In our work, we use the HolisticBias (smith2022imsorryhearthat) metric that measures a form of extrinsic bias. The HolisticBias metric builds on BOLD (Dhamala_2021) and contains a dataset of 566,000 prompts that are categorized into 13 social biases (e.g., Gender, Sexual Orientation, Race and Ethnicity, Nationality, etc.). Bias is quantified by prompting the LLMs with sentences from the different groups and examining the toxicity of the generated text. The toxicity of a text determines how disrespectful, unpleasant, or harmful the language is. The toxicity is computed using a BERT (devlin2019bert) model that is pre-trained on a toxic comment classification dataset [^1] https://www.kaggle.com/c/jigsaw-toxic-comment-classification-challenge. The BERT model outputs a score in $[0,1]$, where a higher score is assigned to more toxic text.
To measure the bias of the LLM, we adopt the procedure outlined in prior works (zayed2024fairness; Dhamala_2021). Each social bias, or group, in the HolisticBias dataset consists of a set of subgroups $G$. For example, the gender group contains subgroups such as binary, queer, transgender, etc. The bias of the LLM $\Theta$ on a set of subgroups $G$ is
(4) $\text{bias}_{\Theta}(G)=\sum_{g\in G}\left|T_{G}-T_{g}\right|$
(5) $\text{where}\quad T_{g}=\frac{1}{|D_{g}|}\sum_{x\in D_{g}}\text{tox}_{\Theta}(x)$
(6) $\text{and}\quad T_{G}=\frac{1}{|G|}\sum_{g\in G}T_{g}$
@@ 3.1. Multi-Head Attention in Transformers
Most modern LLMs for causal language modeling are based on the Transformer architecture proposed by 10.5555/3295222.3295349. On a high level, the Transformer is a neural network that comprises $n$ blocks stacked on top of each other, where each block has a Multi-Head Attention (MHA) layer followed by a fully connected layer. The MHA layer utilizes scaled-dot product attention, which operates on a set of queries $Q$, keys $K$, and values $V$, which come from the previous layer’s output. The query, key, and values are computed from the input text embeddings for the first block. Scaled-dot product attention is computed as
@@ 3.2. Attention Head Pruning for Fairness
Pruning in LLMs refers to the practice of “pruning” or “dropping” parts of the LLM. Pruning strategies can be broadly categorized into two categories: unstructured and structured pruning. Unstructured pruning refers to the practice of pruning at the neuron level, where a subset of neurons are pruned. On the other hand, structured pruning strategies remove entire structural components, such as attention heads or layers, from the LLM. Attention head pruning is a structured pruning strategy that prunes one or more attention heads from multi-head attention in a Transformer block. For example, Figure 1(b) shows multi-head attention with the second and fourth heads pruned. These heads are not concatenated before multiplying with the output projection matrix.
@@
Multiple works have been proposed (DBLP:journals/corr/abs-1905-10650; voita-etal-2019-analyzing; fan2021layerwisemodelpruningbased; han2016deepcompressioncompressingdeep; behnke-heafield-2020-losing) to prune redundant attention heads in LLMs that have a negligible impact on utility. However, they do not account for fairness. We define the problem of pruning attention heads to improve fairness while minimizing loss to model utility as follows
@@
Recently, zayed2024fairness proposed the fairness-aware structured pruning (FASP) algorithm that prunes attention heads while accounting for both fairness and model utility. FASP quantifies the contribution of an attention head by measuring the change in bias and perplexity of LLM before and after pruning it. Formally, for a model with $N$ attention heads, the impact of each head $h$ on a social group represented by set $G$ is estimated as
@@
As the authors concede, FASP fails to account for the complex non-linear relationship of attention heads in an LLM. For example, if dropping the first and second heads decreases bias while not reducing perplexity, dropping both of them together may not decrease bias, and it might worsen perplexity. Nevertheless, FASP is a first step toward fairness-aware attention head pruning.
@@ 4.1. Discrete Optimization with Simulated Annealing
where the set of vertices $V$ consists of all bitstrings of length NN
@@
The cooling schedule determines how the temperature changes over time. In SA, we start with a high temperature and anneal it over time. This encourages exploration in the early stages of the SA search and exploitation towards the end. We adopt the logarithmic cooling schedule (hajek1988cooling). The temperature $T_{i}$ at the $i^{th}$ iteration is defined as
@@
where $T_{0}$ is the initial temperature. We utilize the temperature initialization algorithm proposed by Ben-Ameur2004 to set $T_{0}$.
@@ 5. Experiments
Datasets and Models. We focus on reducing gender bias from the HolisticBias metric in LLMs and consider 6 LLMs for our analysis: Distilgpt-2 (sanh2019distilbert), GPT-2 (radford2019language), GPT-Neo-125M (gpt-neo), GPT-Neo-1.3B (gpt-neo), GPT-J-6B (gpt-j), Llama-2-7B (touvron2023llama2openfoundation). Our choice of social bias and LLMs mimics the state-of-the-art FASP pruning strategy setup for a fair comparison. The LLMs chosen cover a diverse set of parameter sizes, from 82M (Distilgpt-2) all the way to 7B (Llama-2-7B).
The architecture for the surrogate DNNs is shown in Table 2. We find that small DNNs with two hidden layers can effectively predict a state’s bias and perplexity in the search space. The size of the input layer is the number of attention heads in the LLM. The output layer has one neuron. For the smaller LLMs ($1\text{B}$ parameters) with fewer attention heads, we find that hidden layers with more neurons tend to overfit. Thus, we opt to use small hidden layers with 64 and 32 neurons. For the bigger LLMs, hidden layers with 256 and 128 neurons perform better. The size of the datasets used to train the surrogate DNNs is highlighted in Table 2. We use a train-validation split of 95%-5%. The DNNs are trained till the validation split error starts to increase. During the dataset creation in Algorithm 1, we set the upper bound on the number of attention heads dropped $n_{u}$ to up to $10\%-20\%$ of the total attention heads in the LLM. We set $\eta_{bias}\in[0.1,0.2]$ and $\eta_{ppl}=1$. Higher values of $\eta_{bias}$ are chosen for the smaller LLMs since their inference times are faster. Additionally, dropping attention heads increases perplexity and reduces bias since meaningless and random text has a low bias by default. Therefore, we need the perplexity DNN $\theta_{ppl}$ to be as accurate as possible to ensure that the SA search returns states with low bias and low perplexity. To ensure high data quality for the perplexity samples, we used the entire WikiText-2 validation split with $2,461$ text sequences. On the other hand, the HolisticBias validation dataset has $10,268$ gender bias prompts, and we only use a fraction of them to speed up the data collection process. We perform experiments on a shared GPU cluster with variable GPU availability depending on server load. We note that the dataset collection process is offline. Instead of setting a time limit for Algorithm 1, we collect at least 25,000 samples to train the DNNs. Once the surrogate DNNs are trained, they will be used in the online phase for the SA search.
@@
Baseline and Considerations. We compare our approach with four pruning strategies. The first is FASP (zayed2024fairness), which is the state-of-the-art fairness-aware pruning algorithm. The other two strategies prune heads based on weight magnitude (han2016deepcompressioncompressingdeep) and gradient magnitude (DBLP:journals/corr/abs-1905-10650). On a high level, these methods rank the heads based on how large the weights of the heads or gradients are with respect to a loss function after masking the heads. Heads with low weight and gradient magnitude are deemed unimportant. These methods are state-of-the-art general-purpose pruning algorithms. The fourth approach is random pruning, which serves as an ablation study.
@@
where $\mathcal{P}$ is the perplexity dataset, $p_{max}$ is the threshold for maximum perplexity, and $\mathcal{P^{\prime}}$ is the updated dataset after clamping $\mathcal{P}$, and $\sigma$ is the standard deviation threshold. We set $\sigma\leq 10$. After clamping, we scale down the dataset by $p_{max}$ to ensure that all values lie in $[0,1]$. We observe that this approach improves the performance of the DNNs during the search as they assign the value $p_{max}$ to states that degrade the model’s utility beyond acceptable levels.